OpenStudy (anonymous):

Which of the following functions grows the fastest as x grows without bound?

OpenStudy (anonymous):

@acxbox22 @amistre64 @CausticSyndicalist @dan815 @jagr2713 @misty1212 @myininaya @quickstudent @TheSmartOne @triciaal

OpenStudy (16shuston):

so without bound means with out limit. I would replace x with infinity and see which one grows quicker

OpenStudy (anonymous):

@16shuston so how would i set up on of these examples. idk how to solve it

OpenStudy (triciaal):

by elimination it would be A or B has to be an exponential function x is approaching infinity very large so fraction is close to zero

OpenStudy (triciaal):

cos x is limited

OpenStudy (anonymous):

okay so it's e^x then?

OpenStudy (16shuston):

yeah that one is growing the quickest

OpenStudy (anonymous):

can u guys help me with another problem?

OpenStudy (triciaal):

next time open new question

OpenStudy (triciaal):

g(x) = f(2x) function f when x = 0 g (x) = 0 function f when x = 1 g(x) = 2

OpenStudy (anonymous):

@triciaal how do i set this up?

OpenStudy (triciaal):

@amistre64

OpenStudy (amistre64):

youhave a chain rule ...

OpenStudy (amistre64):

g = f(2x) what is g'?

OpenStudy (anonymous):

f(2)+2x

OpenStudy (amistre64):

thats not the chain rule define the chain rule for me

OpenStudy (amistre64):

g and f are the names of functions, they are not variables .. this isnt: g = f, times 2x this is: g composed of x, is equal to f composed of (2x) g(x) = f(2x) , a function within a function requires the chain rule

OpenStudy (anonymous):

@amistre64 chain rule is

OpenStudy (anonymous):